November data: (deaths) Deaths Services Admissions < 48 hours > 48 hours Discharges Medical 250 2...

  1. November data: (deaths)

Deaths

Services

Admissions

< 48 hours

> 48 hours

Discharges

Medical

250

2

5

261

Surgical

105

2

4

103

Pediatric

33

0

1

35

Obstetric

36

1

0

34

Psychiatric

47

1

1

45

Newborn

40

1

1

38

Calculate:

  1. Gross death rate
  2. Net death rate
  3. Newborn death rate
  4. Net death rate for medical service
  5. Clinical service with the lowest gross death rate

please show your steps
Answer #1

a. Gross death rate = (Total number of deaths of inpatients in a period / Total of number of discharges in the same period) * 100

As per given information Total number of deaths of inpatient in a period = Deaths in different services before and after 48 hours  in the month of November =  2+5+2+4+0+1+1+0+1+1+1+1= 19.

Total number of discharges in the same period= Discharges in different services = 261+103+35+34+45+38= 516

  Gross death rate = (19/ 516) * 100 = 3.68 %

Note: For calculate gross death rate, deaths before and after 48 hrs of admission should be include.

b. Net death rate = [Total number of inpatient deaths ( including new born) - deaths < 48 hrs of admission/ total discharges - deaths < 48 hours] * 100

= ( 12 / 509) *100 = 2.36%

c. Newborn death rate = ( Total number of neonatal deaths / Total number of neonatal discharge ) * 100

= ( 2/ 38) * 100 = 5.26 %

d. Net death rate for medical service is calculated by using the same above mentioned formula for net death rate but calculated only for medical service

=  ( 5/ 259) *100 = 1.93%

e. clinical service with the lowest gross death rate is identified by calculating the gross death rate of each services separately by using the same above mentioned formula for gross death rate.

Gross death rate(GDR) for medical services = (7/261) *100 = 2.68%

GDR for surgical services = (6/103) * 100= 5.83%

GDR for paediatric services = (1/ 35) *100 = 2.86%

GDR for obstetric services = ( 1/34) *100 =2.94%

GDR for psychiatric services = (2/ 45) * 100 = 4.44%

GDR for newborn services = (2/48) * 100 = 4.17%

clinical service with the lowest gross death rate is gross death rate for medical services (2.68%)